OpenStudy (anonymous):

y=(secx)^3 find antiderivative

OpenStudy (lgbasallote):

\[\int (\sec x)^3 dx \implies \int \(sec^2 x )(\sec x) dx \implies \int (1+ \tan^2 x) \sec x dx\]

OpenStudy (anonymous):

oh wait lol

OpenStudy (anonymous):

forgot root thnx though heres some medals

OpenStudy (lgbasallote):

haha yeah that gets confusing sometimes :)

OpenStudy (anonymous):

Break apart into sec\(^2\)(x) and sec(x) first Then let u = tan(x) du = sec\(^2\)(x) dx See what to do now when you combine with what @lgbasallote wrote?